Solve each system graphically. Be sure to check your solution. If a system has an infinite number of solutions, use set builder notation to write the solution set. If a system has no solution, state this.
b+ 2a = 2,
a= -3 - b

Answer by mananth(16946)

b+ 2a = 2,
a= -3 - b
equation1
when a= 0 b= 2 ( 0 , 2 )
when b = 0 a= 1 ( 1 , 0 )
Equation 2
when a= 0 b= -3 ( 0 , -3 )
when b = 0 a= -3 ( -3 , 0 )

From graph you find solution is (5,-8)
2 a + 1 b = 2 .............1
1 a + 1 b = -3 .............2

multiply (1)by 1
Multiply (2) by -1
2 a + 1 b = 2
-1 a + -1 b = 3
Add the two equations
1 a = 5
/ 1
a = 5
plug value of a in (1)
2 a + 1 b = 2
10 + 1 b = 2
1 b = 2 -10
1 b = -8
b = -8.00
m.ananth@hotmail.ca